OpenStudy (anonymous):

PHYSICS HELP (Show me how to do work as well) A 1.0 x 10^4 kg spacecraft is traveling through space with a speed of 1200 m/s relative to Earth. A thruster fires for 2.0 mins, excerting a continious force of 25kN on the spacecraft in a direction opposite the spacecraft's motion. Calculate the initial momentum and final momentum of the space craft?

OpenStudy (loser66):

I know initial momentum only, hihihi... because it 's easy p = mv . that's it. just plug the number in

OpenStudy (loser66):

oh, now I know how to calculate the final momentum . the momentum of the system after collision is \(M_{thruster} + M_{scraft}\) now calculate the momentum of the thruster. it is F = \(\dfrac{dP}{dt}\) so, P = F t = 25kN *0.2 (min) = 3* 10^5 kg m/s because it is opposite direction to the spacecraft, so it is has minus sign in the front so, final momentum of the craft is mv - 3*10^5 = 1.0*10^4 * 1200 - 3*10^5 = 117 *10^5 kg m/s

OpenStudy (anonymous):

\[3 x 10^{5} \] is the initial momentum of the spacecraft right?

OpenStudy (loser66):

nope, initial momentum of the spacecraft is 1200* 10^4 = 1.2 *10^7 kg m/s 3 * 10^5 is momentum of the thruster